EXPERIMENTAL MAPPING OF THE ABSOLUTE VALUE OF THE ELECTRONIC TRANSITION DIPOLE MOMENT FUNCTION μ_{e}(R) OF THE Na_{2} A^{1}Σ_{u}^{+} − X^{1}Σ_{g}^{+} SYSTEM

\underline{E. AHMED}, B. BESER, P. QI, S. KOTOCHIGOVA, Physics Department, Temple University, Philadelphia, PA 19122; J. HUENNEKENS, Physics Department, Lehigh University, Bethlehem, PA 18015; A. M. LYYRA, Physics Department, Temple University, Philadelphia, PA 19122.

The absolute value of the transition dipole moment function μ_{e}(R) of the A^{1}Σ_{u}^{+} − X^{1}Σ_{g}^{+} band of Na_{2} was mapped experimentally within a relatively large range of internuclear distance R. The transition dipole moment matrix element of a set of rovibrational transitions between the A^{1}Σ_{u}^{+} and X^{1}Σ_{g}^{+} states was measured using the Autler-Townes (AT) effect^{a}. By employing the R-centroid approximation, or a fit with a polynomial function involving the higher R-moments, μ_{e} as function of the internuclear distance was constructed. The AT effect based measurements give the absolute value of μ_{e} which can be used to ”normalize” the transition dipole moment functions μ_{e}(R) obtained with other experimental techniques, such as line intensity based measurements, which in general give only the relative behavior of μ_{e}(R).

aE. Ahmed et al. J. Chem. Phys. 124, 084308 (2006)